If 5.5 g of sodium chloride reacts with 2.3 g of magnesium bromide, what is the limiting reactant



1
Expert's answer
2021-03-16T08:25:58-0400

moles of NaCl = 5.5/58.44 = 0.0941mol

moles of MgBr2 = 2.3/184.113 = 0.0124mol

MgBr2 is a limiting reactant